# Why do financial prices exhibit Brownian motion despite predictable order flow?

### v3

@@ Abstract
In financial market microstructure, there are two enigmatic empirical laws: (i) the market-order flow has predictable persistence due to metaorder splitters by institutional investors, well formulated as the Lillo-Mike-Farmer model. However, this phenomenon seems paradoxical given the diffusive and unpredictable price dynamics; (ii) the price impact $I(Q)$ of a large metaorder $Q$ follows the square-root law, $I(Q)\propto\sqrt{Q}$. Here we theoretically reveal why price dynamics follows Brownian motion despite predictable order flow by unifying these enigmas. We generalize the Lillo-Mike-Farmer model to nonlinear price-impact dynamics, which is mapped to an exactly solvable Lévy-walk model. Our exact solution shows that the price dynamics remains diffusive under the square-root law, even under persistent order flow. This work illustrates the crucial role of the square-root law in mitigating large price movements by large metaorders, thereby leading to the Brownian price dynamics, consistently with the efficient market hypothesis over long timescales.
@@ Introduction.
There are two empirical enigmas in financial market microstructure [1, 2, 3]. The first enigma is the long memory of the market-order flow. A market order is an immediate decision to buy or sell stocks, which is represented by $\epsilon_{t}=+1$ ($\epsilon_{t}=-1$) for the buy (sell) market order at time $t$ (e.g., $\{\epsilon_{t}\}_{t\geq 1}=\{+,+,+,-,+\}$ signifies three buys, one sell, and one buy orders). The long memory implies that such order signs are easily predictable and, thus, the autocorrelation (ACF) of the market-order signs decays very slowly:
@@
Here we propose a minimal model that naturally extends the LMF framework to incorporate nonlinear price-impact dynamics, thereby unifying these two enigmas from a single, coherent perspective. We consider the system composed of $M$ order splitters whose metaorder-size statistics follow the power law $\psi_{m}(Q)\propto Q^{-\alpha-1}$ with $21<\alpha<$ 2. All traders are assumed to execute their metaorders whose price impact obeys the nonlinear scaling $I(Q)\propto Q^{\delta}$ with $0<\delta\leq 1$. Crucially, this model is exactly solvable. The price-impact contribution by a single trader can be mapped to the Lévy-walk theory [15, 16] with nonlinear walking-speed down [17]. Our exact solution shows that the price dynamics is always diffusive even in the presence of the long memory if $\delta=1/2$. This result implies that the SRL plays the crucial role in mitigating the price impact due to the huge liquidity consumption by institutional investors. Additionally, we find that this model exhibits consistent behavior with other empirical laws (such as the inverse-cubic law [1, 3, 18, 19, 20, 21, 22] and volatility clustering [3, 2]). Thus, our simple model is very simple but minimally consistent with various empirical enigmas in finance with clear exact solutions.
@@ I.1 Appendix A: the Lévy-walk theory
Let us consider a single particle obeying Lévy walks with the flight-time PDF $\psi_{m}(t)=\alpha t^{-\alpha-1}\Theta(t-1)$, the corresponding flight distance $\Delta x=t^{\delta}$, and the exponential resting-time distribution $\psi_{r}(t)=(1/\tau_{r})e^{-t/\tau_{r}}$. The space-time coupling function is given by $\psi(x,t):=(1/2)\delta(|x|-t^{\delta})\psi_{m}(t)$. The PDF of the particle displacement $x$ at time $t$ is written as $P(x,t)$, and its Fourier-Laplace representation is written as $P(k,s):=\int_{0}^{\infty}{\rm d}te^{-ts}\int_{-\infty}^{\infty}{\rm d}xe^{-ts-{\rm i}kx}P(x,t)$. In the following, the arguments $k$ and $s$ always signify the Fourier-Laplace representation.
@@ I.3 Appendix C: the inverse-cubic law
Figure 4: (a) Complementary cumulative distribution function for the price changes, showing the inverse-cubic law $P_{>}(\Delta m)\propto(\Delta m)^{-\beta}$ with $\beta=3$. We assumed the following parameters: $\delta=1/2$, $\alpha=1.5$, $\tau_{r}=10^{5}$, $t=10^{4}$, and $M=1$. See the inset for the corresponding semi-log plot. (b) Volatility ACF in our model, numerically showing the long memory $C_{V}(\tau)\propto\tau^{-\zeta}$ with $\zeta\in(0,1)$ for the parameters $M=1$, $\delta=0.5$, $\tau_{r}=10$, and $t=1000$. Empirically, the power-law exponent obeys $\zeta\approx\alpha-1$. More detailed numerical investigation is presented in the SI.

### v4

@@ Abstract
In econophysics of financial market microstructure, there are two enigmatic empirical laws: (i) the market-order flow has predictable persistence due to metaorder splitters by institutional investors, well formulated as the Lillo-Mike-Farmer model. However, this phenomenon seems paradoxical given the diffusive and unpredictable price dynamics; (ii) the price impact $I(Q)$ of a large metaorder $Q$ follows the square-root law, $I(Q)\propto\sqrt{Q}$. Here we theoretically reveal why price dynamics follows Brownian motion despite predictable order flow by unifying these enigmas. We generalise the Lillo-Mike-Farmer model to nonlinear price-impact dynamics, which is mapped to an exactly solvable Lévy-walk model. Our exact solution shows that the price dynamics remains diffusive under the square-root law, even under persistent order flow. This work illustrates the crucial role of the square-root law in mitigating large price movements by large metaorders, thereby leading to the Brownian price dynamics, consistent with the efficient market hypothesis over long timescales.
@@ Introduction.
There are two empirical enigmas in econophysics [1, 2, 3]. The first enigma is the long memory of the market-order flow. A market order is an immediate decision to buy or sell stocks, which is represented by $\epsilon_{t}=+1$ ($\epsilon_{t}=-1$) for the buy (sell) market order at time $t$ (e.g., $\{\epsilon_{t}\}_{t\geq 1}=\{+,+,+,-,+\}$ signifies three buys, one sell, and one buy orders). The long memory implies that such order signs are easily predictable and, thus, the autocorrelation (ACF) of the market-order signs decays very slowly:
@@
Here we propose a minimal model that naturally extends the LMF framework to incorporate nonlinear price-impact dynamics, thereby unifying these two enigmas from a single, coherent perspective. We consider the system composed of $M$ order splitters whose metaorder-size statistics follow the power law $\psi_{m}(Q)\propto Q^{-\alpha-1}$ with $21<\alpha<$ 2. All traders are assumed to execute their metaorders whose price impact obeys the nonlinear scaling $I(Q)\propto Q^{\delta}$ with $0<\delta\leq 1$. Crucially, this model is exactly solvable. The price-impact contribution by a single trader can be mapped to the Lévy-walk theory [15, 16] with nonlinear walking-speed down [17]. Our exact solution shows that the price dynamics is always diffusive even in the presence of the long memory if $\delta=1/2$. This result implies that the SRL plays the crucial role in mitigating the price impact due to the huge liquidity consumption by institutional investors. Additionally, we find that this model exhibits consistent behavior with other empirical laws (such as the inverse-cubic law [1, 3, 18, 19, 20, 21, 22] and volatility clustering [3, 2]). Thus, our simple model is very simple but minimally consistent with various empirical enigmas in econophysics of market microstructure with clear exact solutions.
@@ I.1 Appendix A: the Lévy-walk theory
Let us consider a single particle obeying Lévy walks with the flight-time PDF $\psi_{m}(t)=\alpha t^{-\alpha-1}\Theta(t-1)$, the corresponding flight distance $\Delta x=t^{\delta}$, and the exponential resting-time distribution $\psi_{r}(t)=(1/\tau_{r})e^{-t/\tau_{r}}$. The space-time coupling function is given by $\psi(x,t):=(1/2)\delta(|x|-t^{\delta})\psi_{m}(t)$. The PDF of the particle displacement $x$ at time $t$ is written as $P(x,t)$, and its Fourier-Laplace representation is written as $P(k,s):=\int_{0}^{\infty}{\rm d}t\int_{-\infty}^{\infty}{\rm d}xe^{-ts-{\rm i}kx}P(x,t)$. In the following, the arguments $k$ and $s$ always signify the Fourier-Laplace representation.
@@ I.3 Appendix C: the inverse-cubic law
Figure 4: (a) Complementary cumulative distribution function for the price changes, showing the inverse-cubic law $P_{>}(\Delta m)\propto(\Delta m)^{-\beta}$ with $\beta=3$. We assumed the following parameters: $\delta=1/2$, $\alpha=1.5$, $\tau_{r}=10^{5}$, $t=10^{4}$, and $M=1$. See the inset for the corresponding semi-log plot. (b) Volatility ACF in our model, numerically showing the long memory $C_{V}(\tau)\propto\tau^{-\zeta}$ with $\zeta\in(0,1)$ for the parameters $M=1$, $\delta=0.5$, $\tau_{r}=1$, and $t=10$. Empirically, the power-law exponent obeys $\zeta\approx\alpha-1$. More detailed numerical investigation is presented in the SI.